If the roots of x³+px²+qx+r=0 are in G.P., then find the relation between p, q, r.

Grade:11

1 Answers

Dilli PRASAD
6 years ago
Let a/r, a, ar be the roots (roots are in gp)Product of roots a/r*a*ar = -ra^3=-r that implies a=-(r)^1/3Put a value in given eqtsWe get p^3r+q^3=0